one angle is 30 degrees more than twice the size of another angle.
If the angles are complementary, how many degrees are in each angle?

Answer by checkley77(12844)

X+30+2X=90
3X=90-30
3X=60
X=60/3
X=20 DEGREES FOR THE SMALLER ANGLE.
90-20=70 DEGREES FOR THE LARGER ANGLE.
PROOF:
20+30+2*20=90
20+30+40=90
90=90